Thursday, January 13, 2022
We are having connectivity issues when the end user is connecting to our network using Airtel, BSNL, Tata India. We have contacted them and waiting for a reply. Your VPS's are all up and working when connected from all other network providers. Unfortunately, there is nothing we can do to resolve the issue as it's not caused by us. Many other hosting providers are having the exact same...
Tuesday, December 7, 2021
One of our payment processors, Paddle is having technical difficulties for the past few hours. This is causing payments not getting completed or payment screen is not opening. They have notified us that they are working on resolving the issue. If your payment is urgent, please use other methods.
Tuesday, November 9, 2021
We are performing an emergency maintenance for lavz7highram3 due to software stability issues. All VPS's are getting restarted due to this. Update: 01:36 AM CST: All VPS's are back online.
Thursday, September 30, 2021
T hursday, 30 September 2021, 12:17 - We are investigating ukstorage1 connectivity issues and will post more information here as we have it. Thursday, 30 September 2021, 12:49 - The server has a hardware problem. We were able to bring it back online but we need to schedule a downtime to replace the faulty memory module. Friday, 01 October 2021, 03:42 - The server has become unstable again...
Tuesday, August 10, 2021
T uesday, 10 August 2021, 16:02 - we're investigating connectivity issues with 104.237.11.x in Los Angeles. A ticket has been opened with the datacenter. We will post more information here as we have it. If your IP starts with 104.237.11 and your VPS is currently inaccessible, it's due to this situation. Tuesday, 10 August 2021, 16:14 - connectivity has been restored, we're still waiting for...
Tuesday, August 10, 2021
T uesday, 10 August 2021, 08:50 - we're investigating an issue with connectivity to the 63.142.243.xxx range in our Los Angeles hosting. We'll post updates here as we have them. Tuesday, 10 August 2021, 09:04 - looks like an issue with the datacenter. We have opened a ticket with them. We'll post updates here as we have them.
Tuesday, June 22, 2021
We are shutting uklegvz7highram5 down to replace a faulty memory stick which is causing instability issues. The DC techs will replace it and put the server back on as soon as possible. Update 10:45: This has been completed.
Thursday, May 27, 2021
Thursday, 27 May 2021, 21:40 We are currently experiencing problems with routing in our Dallas Network. Our engineers are working on it. Thursday, 27 May 2021, 22:37 Service should now be restored to the affected ISPs/carriers. We're still waiting for a final confirmation from the vendor and an RFO.
Wednesday, April 14, 2021
Wednesday, 14 April 2021, 13:55 - we're investigating connectivity issues with our lavz7highram2 node now. We'll update this announcement as we have more info. Wednesday, 14 April 2021, 14:35 - this has been resolved and all VMs are back online.
Tuesday, April 6, 2021
Tuesday, 06 April 2021, 09:52 - we're investigating connectivity issues with our nj1vz7highram1 server now. We will update this announcement as we have more information. Tuesday, 06 April 2021, 10:07 - server appears to be back, awaiting confirmation of stability from the datacenter.